OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x + 2y = -6 3x + y = 8

OpenStudy (precal):

-x+2y=-6 -x=-2y-6 x=2y+6

OpenStudy (anonymous):

Lifesaver :)

OpenStudy (precal):

Do you realize there are many ways to do this system by this method?

OpenStudy (anonymous):

No I didn't know that, I haven't learned this yet at all. I'm taking some sort of "pretest". and I got -2y + 6 somehow

OpenStudy (precal):

well do not test out of systems, it is important to learn all of the different methods. This will help you in the future math courses.

OpenStudy (anonymous):

Thank you :)

OpenStudy (precal):

anytime :)